OpenStudy (anonymous):

Annika Sorenstam won the 2006 MasterCard Classic with a final score of 8 under par, or -8. Her scores for the first two of the three rounds were -5 and -1. What was Ms. Sorenstam's score for the third round?

OpenStudy (texaschic101):

-5 + (-1) + x = -8 -5 - 1 + x = -8 -6 + x = -8 x = -8 + 6 x = -2